Find the general solution except when the exercise stipulates otherwise:

1. a) (D3 + 7D2+ 19D + 13) y =0; whenx=0, y=0, y'=2, y''=-12

b) (D3 + D2 + 4D +4) y =0; when x=0, y=0, y'=-1, y''=5

c) d2x/dt2 + 2b (dx/dt) +k2x = 0, k>b>0; when t=0, x=0, dx/dt = v0
Answer #1

org y = 2 when on=0 2 = - ¢, 4 (263 – 3 C) G + 3 cq - 203 - 2 Again yo = e, ed + e 14% (66, +463) Sin 24 = (460+603)assa 46 cG diffe given general eqr be (03 +54 AD 1479 = 0 eman be a trial son of the eam in + m + 4m + 4 = 0 A.E. m 3 smo bm+1) +4(mar2 Cz - 1 C2 = 0 = 7C3=0 ( el general 60th be y = -x +4 Co82 +2b at2 Let A.E., m² 2bm given diffen eqr. be dx dale dz tkon=0 k